OpenStudy (anonymous):

Solve the proportion (-5)/(x-4) = 1/(x+7)

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

(-5)(x+7) = (x-4) -5x - 35 = x-4 -31 = 6x x = -31/6

OpenStudy (anonymous):

right